# Simple Methods Defend RAG Systems Well Against Real-World Attacks

### v1

@@ Methods
In the first study, we evaluated OOD detection on four query-document datasets spanning 16 domains. For each domain, domain-specific queries served as positive examples, and an equal number of queries from other domains were used as negatives. We split data 90:10 for training and testing with balanced classes. Additional experiments used COVID-19 samples as positives and 4chan or LLM-generated queries as negatives. Embeddings were generated with all-mpnet-base-v2, and we set $k=200$ PCs for both EVR and p-values criteria. [^4] See Appendix for $k$ and model ablations. Semantic search algorithms tuned the radius and $m$ values, while ML models tuned $m$ only.
For the NC-based method, we used $\lambda=0.5,\alpha=1.0,\beta=1.0$. CLS embeddings from bert-base-uncased were clustered (7 clusters for COVID-19, 3 for SU), and fine-tuned for 3 epochs with batch size 16 and learning rate $2\times 10^{-5}$.
@@
In the second study, we evaluated end-to-end RAG. We retrieved the top-10 similar queries using all-MiniLM-L12-v2, re-ranked them with cross-encoder/ms-marco-MiniLM-L-6-v2, and using the top3, we generated responses with GPT-4o (GPT-3.5-turbo for SU). Ten annotators each rated relevance and correctness on a 5-point Likert scale; each sample was rated by two annotators (600 total annotations). Annotators with $<$0.20 average Cohen’s kappa were excluded and reannotations were collected. LLM-as-a-Judge was GPT-4o.
EVR criterion p-values criterion Dataset $\epsilon$-ball $\epsilon$-cube $\epsilon$-rect LogReg SVM GMM $\epsilon$-ball $\epsilon$-cube $\epsilon$-rect LogReg SVM GMM COVID-19 0.942 0.937 0.850 0.981 0.985 0.937 0.985 0.951 0.942 0.966 0.976 0.942 Substance Use 0.940 0.928 0.936 0.967 0.972 0.964 0.944 0.944 0.904 0.964 0.976 0.960 Stackexchange 0.910 0.904 0.885 0.963 0.962 0.942 0.913 0.910 0.810 0.961 0.963 0.949 MSMARCO 0.896 0.890 0.869 0.925 0.930 0.891 0.876 0.869 0.809 0.933 0.935 0.917 Average #PCs 7 6 4 140 143 40 13 15 5 125 136 65
Table 1: The results for both EVR and p-values criteria across 4 datasets, along with the average number of PCs used. Best result per dataset per criterion in bold, second best underlined.
Results
Out-Of-Domain Detection
Table 1 shows all datasets results. For each dataset-method combination, a hyperparameter search was performed. The average number of PCs used for each method is reported in the last row (see Appendix for unaggregated results).
Furthermore, we evaluated our methods under practical conditions where some ID data and auxiliary OOD data generated by LLMs are available. For the COVID-19 domain, we considered two settings: (1) an isolated test set containing both COVID-19 and LLM-attack queries, and (2) the 4chan dataset, which reflects a real-world scenario. For the SU domain, we evaluated only on an isolated set comprising SU and LLM-attack samples (results shown in Table 2).
For the NC technique (Figure 3), we visualized the learned feature space projected into three dimensions: the two most dominant class weight directions ($w_{1}$, $w_{2}$) and the principal component ($u_{3}$) capturing OOD variance. The selected classes correspond to the two most represented clusters in each dataset, as obtained through our earlier clustering step. The “Outlier” points correspond to unseen OOD queries (see Appendix for all cases).
(a) COVID-19 dataset.
(b) Substance Use dataset.
Figure 3: Visualization of features projected into the three-dimensional space consisted of $w_{1},w_{2}$ (the weights of the last fully connected layer of the respective labeled classes) and the principal eigenvector of OOD features. The selected Classes are the ID classes with the highest representation for each dataset, and the Outlier means features of test unseen OOD data. We observe that the feature separability between ID and OOD is clearer in this 3D space.
EVR criterion p-values criterion NC Dataset $\epsilon$-ball $\epsilon$-cube $\epsilon$-rect LogReg SVM GMM $\epsilon$-ball $\epsilon$-cube $\epsilon$-rect LogReg SVM GMM C19 LLM-Attack 0.937 (79.1) 0.937 (44.2) 0.937 (44.2) 0.903 0.903 0.806 0.937 (43.7) 0.922 (77.2) 0.942 (53.4) 0.864 0.893 0.791 0.940 C19 4chan 0.682 (23.4) 0.766 (0.0) 0.766 (0.0) 0.582 0.597 0.607 0.761 (0.5) 0.682 (27.3) 0.771 (20.4) 0.572 0.607 0.512 0.622 SU LLM-Attack 0.729 (100.0) 0.760 (100.0) 0.708 (96.9) 0.922 0.932 0.594 0.870 (92.7) 0.859 (99.5) 0.849 (99.5) 0.932 0.938 0.875 0.906
Table 2: Accuracy (%non-empty) across datasets and methods, where %non-empty is the percentage of questions with at least one neighbor within the border. COVID-19 (C19), Substance-Use (SU), Neural Collapse (NC).
RAG Evaluation
@@ Results
Table 3 shows results of the annotation task and LLM-as-a-Judge for both COVID-19 and SU domains (see Appendix for unaggregated results and further error analysis). For each dimension, we conducted independent t-tests to determine whether there were significant differences between ID and OOD responses. We report the mean and standard deviation of Relevance and Correctness scores for each group, along with their corresponding $p$-value.
@@
Table 3: Mean ($\pm$±Standard Deviation) of both dimensions for the different groups of in-domain (ID) and OOD (OOD) questions. C19 denotes the COVID-19 domain. SU denotes the Substance Use domain. 18 cases were marked as ”N/A” for Correctness, as it is not possible to assess them scientifically.
In Table 4, we present the evaluation of 300 retrieved samples using our best-performing model, a GMM. The model was trained with COVID-19 samples as positive instances and LLM-attack samples as negative instances, as a practitioner would do. Additionally, we report the performance of GPT-4o, prompted to function as an OOD detector. We experimented with three variations: (1) providing GPT-4o with all positive samples to establish a KB, (2) supplying both positive (COVID-19) and negative (LLM-attack) samples, and (3) using a 20-shot setup with examples from both classes. The table only shows results for the best-performing approach (variation 2). For the prompts and more prompt optimization results, please refer to Appendix.
@@
Table 4: GMM and GPT-4o results in the dataset of 150 in-domain (ID) and 150 out-of-domain (OOD) samples. We report the number of True Positives (TP) and False Negatives (FN) for each category, along with the average relevance and correctness scores.
Discussion
Out-Of-Domain Detection
Our findings indicate that across multiple datasets, the two criteria developed—EVR and p-value —do not lead to significant differences in classification accuracy (Table 1). While in the COVID-19 domain, the p-value criterion improves the performance of semantic search algorithms in the COVID-19 domain, it does not improve machine learning algorithms. Similarly, for most datasets, except for MSMARCO, the p-value criterion enhances results for the $\epsilon$-ball and $\epsilon$-cube algorithms.
An important observation is that the p-value criterion tends to select a larger number of PCs for semantic search algorithms. This is because the new ranking approach prioritizes more relevant PCs, allowing their early inclusion and effective combination with others to enhance performance. In contrast, machine learning algorithms do not exhibit the same effect, possibly due to the higher number of selected PCs.
Moreover, although semantic search algorithms generally have lower accuracy compared to machine learning approaches, their performance remains relatively similar (with the exception of $\epsilon$-rect). This suggests that these methods are proper alternatives, especially given their ability to operate with a significantly smaller number of PCs. In this way, we can reduce complexity and thus enhance interpretability and computational efficiency, making them attractive for applications where error case analysis is critical, such as the medical domain.
We further conducted a qualitative analysis to uncover patterns in the PCs that were most consistently prioritized by the p-values criterion. Unlike traditional dimensionality reduction criteria, this ranking approach highlights PCs that align closely with meaningful and interpretable themes within each domain. These include, for instance, vaccine eligibility concerns and risk perception in the COVID-19 dataset, or treatment-related and concealment strategies in the SU dataset (see Appendix). Our analysis reveals that PCs selected via the p-values criterion are not only discriminative for classification, but also interpretable, offering valuable insights into the underlying user concerns.
For the NC technique (Figure 3), the ID classes are highly compact and aligned with the respective weight axes, consistent with this phenomenon, where the penultimate features of ID samples within a class are nearly identical to the last FC layer’s weights of the corresponding class. Meanwhile, OOD queries are clearly separated along the orthogonal OOD component axis ($u_{3}$), showing the effectiveness of our feature separation approach.
In Table 2, the NC method achieves the highest accuracy for the COVID-19 domain, outperforming all machine learning classifiers. While some semantic search algorithms reach comparable performance, NC provides a more stable and interpretable alternative, as the percentage of non-empty cases (questions with at least one neighbor within borders) is small, especially for the 4chan dataset. In the SU domain, the machine learning models seem to perform relatively better, especially for the p-values criterion, with the NC method having a competitive performance of over 0.9 accuracy.
RAG Evaluation
In our evaluation of RAG, we observed that toxic (OOD) samples were able to bypass LLM guardrails (see Appendix). Since RAG systems operate over a fixed knowledge base (KB), it is crucial to ensure that generated responses remain both relevant and accurate. Results from both human annotators and the LLM-as-a-Judge indicate that OOD questions significantly reduce response relevance (Table 3), though no significant difference was found in Correctness scores between ID and OOD queries.
Further insights from Table 4 show that our best model performs comparably to the strongest GPT-4o-based setup. While GPT-4o still achieves higher accuracy (for further prompt optimization, see Appendix), our approach offers similar performance with significantly lower cost, latency, and greater interpretability.
Interestingly, for ID queries, those correctly classified as ID received higher relevance and correctness scores than those misclassified as OOD. Conversely, OOD queries misclassified as ID also showed higher quality scores than correctly rejected OODs. This suggests that even in cases of misclassification, the model tends to favor safer, more useful responses, thereby reducing risk.
In the SU domain, ID queries also scored significantly higher in relevance than OOD ones, though both groups maintained high correctness scores (Table 3). This likely reflects the model’s adherence to prompts, with low relevance attributed to unanswerable or poorly retrieved context. Results from GPT-4o corroborate this trend (see Appendix).
Conclusion
In this study, we introduce a simple yet effective approach for OOD detection in RAG systems by applying PCA to document embeddings and projecting query embeddings onto the resulting subspace. We evaluate two ranking strategies for PC selection: EVR and a p-value-based criterion that prioritizes components best separating ID and OOD queries. In parallel, we adapt an NC-based feature separation method, incorporating clustering and fine-tuning to enhance ID–OOD separability. This approach shows strong performance, particularly in domains like COVID-19, highlighting its value as a high-capacity alternative. We assess three semantic-space algorithms and three machine learning classifiers across four datasets and 16 domains, demonstrating reliable ID vs OOD discrimination. While both PC ranking criteria yield similar results, the p-value method selects more informative components for semantic search and achieves better performance in key domains. Compared to GPT-4o-based OOD detection, our models achieve comparable effectiveness with significantly lower complexity, cost, and improved interpretability.
Beyond classification, we examine real-world implications by assessing how OOD detection impacts the Relevance and Correctness of RAG system responses. Human and LLM-based evaluations confirm a significant drop in relevance for OOD queries, though correctness remains largely unaffected. However, we show that when our method classifies certain queries as ID, even if they are technically OOD, they actually yield higher relevance and correctness.
Our findings suggest that simple, low-dimensional representations can enable efficient, cost-effective, and interpretable OOD detection, as an alternative to modern LLMs.

### v2

@@ 4.2 Experimental Setup
In the first study, we evaluated OOD detection on four query-document datasets spanning 16 domains. For each domain, domain-specific queries served as ID examples, and an equal number of queries from other domains were used as OOD. We split data 90:10 for training and testing with balanced classes. Additional experiments used COVID-19 samples as ID and 4chan or LLM-generated queries as OOD. Embeddings were generated with all-mpnet-base-v2, and we set $k=200$ PCs for both EVR and p-values criteria. [^3] See Appendix C.1 for $k$ and model ablations. Semantic search algorithms tuned the radius and $m$ values, while ML models tuned $m$ only.
@@
In the second study, we evaluated end-to-end RAG. We retrieved the top-10 similar queries using all-MiniLM-L12-v2, re-ranked them with cross-encoder/ms-marco-MiniLM-L-6-v2, and using the top3, we generated responses with GPT-4o (GPT-3.5-turbo for SU). Ten annotators each rated relevance and correctness on a 5-point Likert scale; each sample was rated by two annotators (600 total annotations). Annotators with <<0.20 average Cohen’s kappa were excluded and reannotations were collected. LLM-as-a-Judge was GPT-4o. We refer to Appendix C.1 for more details regarding the annotation process.
Method COVID-19 Subst. Use StackEx MSMARCO C19 LLM-Att C19 4chan SU LLM-Att m
EVR criterion
$\epsilon$-ball 0.942 0.940 0.910 0.896 0.937 0.682 0.729 7
$\epsilon$-cube 0.937 0.928 0.904 0.890 0.937 0.766 0.760 6
$\epsilon$-rect 0.850 0.936 0.885 0.869 0.937 0.766 0.708 4
LogReg 0.981 0.967 0.963 0.925 0.903 0.582 0.922 140
SVM 0.985 0.972 0.962 0.930 0.903 0.597 0.932 143
GMM 0.937 0.964 0.942 0.891 0.806 0.607 0.594 40
p-values criterion
$\epsilon$-ball 0.985 0.944 0.913 0.876 0.937 0.761 0.870 13
$\epsilon$-cube 0.951 0.944 0.910 0.869 0.922 0.682 0.859 15
$\epsilon$-rect 0.942 0.904 0.810 0.809 0.942 0.771 0.849 5
LogReg 0.966 0.964 0.961 0.933 0.864 0.572 0.932 125
SVM 0.976 0.976 0.963 0.935 0.893 0.607 0.938 136
GMM 0.942 0.960 0.949 0.917 0.791 0.512 0.875 65
Baselines
Mahalanobis 0.990 0.800 0.945 0.921 0.944 0.761 0.662 768
kNN 0.976 0.976 0.958 0.935 0.743 0.706 0.771 768
ViM 0.937 0.808 0.874 0.813 0.937 0.766 0.708 768
ODIN 0.966 0.976 0.963 0.945 0.849 0.353 0.906 768
Table 1: OOD detection accuracy across clean datasets (left block) and attack datasets (right block). Methods are grouped by feature-selection criterion (EVR vs p-values) and compared against strong embedding-level baselines. Best per group in bold, second-best underlined. The last column (m) is the average number of dimensions (or PCs) that are used in each method.
5 Analysis
5.1 Out-Of-Domain Detection
Baselines
We compare our approach with four standard baselines from the OOD literature.
Mahalanobis. Following Lee et al. (2018), we fit a Gaussian distribution to in-domain query embeddings and use the Mahalanobis distance to the estimated mean as an OOD score. Queries that lie far from the in-domain embedding manifold are classified as out-of-domain.
ViM. Virtual Logit Matching (ViM) (Wang et al., 2022) combines the confidence of the classifier with the geometry of the feature-space by measuring the residual norm of a query outside the principal subspace of in-domain representations and adjusting it using the classifier logits. Higher scores indicate a greater likelihood of being out-of-domain. Logistic Regression is used as a classifier for logits.
kNN. The k-nearest neighbors (kNN) baseline detects OOD queries based on their distance to the nearest in-domain examples in embedding space (Sun et al., 2022). Queries whose average or $k$-th nearest neighbor distance exceeds a threshold are considered out-of-domain. In our runs, $k=3$.
ODIN. ODIN (Liang et al., 2017) enhances maximum-softmax-based OOD detection by combining temperature scaling with a small, gradient-based input perturbation that amplifies the separation between ID and OOD confidence scores.
The left part of Table 1 reports OOD detection accuracy across four datasets; the final column shows the average number of retained PCs or dimensions for each method (unaggregated results appear in Appendix C.1). Across datasets, the two feature-selection criteria achieve comparable performance, indicating that both reliably identify informative subspaces for OOD detection. However, systematic differences emerge across method families. For semantic-search methods ($\epsilon$-ball and $\epsilon$-cube), the p-value criterion consistently matches or outperforms EVR on three of the four datasets, with the largest gains in the COVID-19 domain. This reflects the p-value ranking’s emphasis on PCs that maximize ID–OOD separability rather than overall variance.
A key consequence of the p-value ranking is that it selects slightly more PCs for semantic-search methods by prioritizing statistically informative dimensions, while remaining highly compact (typically fewer than 15 PCs). In contrast, supervised classifiers (LogReg, SVM, and GMM) are largely insensitive to the ranking criterion, as they operate in much higher-dimensional regimes where PC ordering has limited effect.
Although semantic-search methods slightly trail supervised classifiers, they remain competitive, particularly $\epsilon$-ball and $\epsilon$-cube. Among embedding-level baselines, kNN, which is the closest to our methods, achieves comparable accuracy on some datasets but operates in the full 768-dimensional space. By contrast, our methods use two orders of magnitude fewer dimensions, yielding a favorable trade-off between accuracy, computational efficiency, and interpretability. ODIN performs best on public datasets and remains competitive on real-world ones.
5.2 Attacks
We further evaluate our methods under more practical conditions in which both ID data and auxiliary OOD data generated by LLMs are available. The same set of baselines is used for comparison.
For the COVID-19 domain, we consider two complementary evaluation settings. First, we construct an isolated test set containing a mixture of COVID-19 and LLM-generated attack queries. Second, we evaluate on the 4chan dataset, which represents a realistic scenario with naturally occurring, noisy, and adversarial user queries. For the SU domain, we evaluate on an isolated test set comprising SU and LLM-attack queries. We report results in the right part of Table 1.
Overall, results under attack settings largely mirror those on clean data (Section 5.1), but with reduced robustness. Semantic-search methods remain competitive in accuracy, though the learned semantic space contracts substantially, causing most test queries to fall outside its boundaries. The advantage of supervised methods narrows in harder settings, such as LLM-generated attacks, where distributional shift and noise are less pronounced.
Similar trends hold for embedding-level baselines: while they can achieve competitive accuracy, they operate in the full embedding space and incur substantially higher dimensional and computational cost. Notably, although ODIN performs well on clean data, its accuracy degrades sharply under attack (particularly on the 4chan dataset) likely because its reliance on confidence amplification via temperature scaling and small input perturbations is ineffective when adversarial or noisy OOD queries remain confidently mapped to ID regions.
5.3 RAG Evaluation
@@
Table 2 reports results from the annotation study and the LLM-as-a-Judge evaluation for the COVID-19 and SU domains; unaggregated results and additional error analysis appear in Appendix C.2. For each dimension, we conduct independent $t$-tests comparing ID and OOD responses, reporting mean, standard deviation, and corresponding $p$-values for Relevance and Correctness.
Across both domains, OOD queries result in a statistically significant reduction in response relevance, consistently observed by human annotators and the LLM-as-a-Judge. In contrast, correctness scores do not differ significantly between ID and OOD responses. This suggests that while RAG systems often remain factually correct under OOD inputs, they frequently fail to produce responses aligned with user intent. We further find that toxic or adversarial OOD queries can bypass LLM guardrails (Appendix C.2), underscoring the need for effective external OOD detection in RAG pipelines.
Next, we evaluate the downstream performance of a full RAG system that relies on LLMs as standalone OOD detectors against a two-stage pipeline in which our method is used as an external OOD detection module. Table 3 reports a comparative evaluation on 300 retrieved samples using our best-performing model (GMM) and GPT-4o configured for OOD detection, where the GMM is trained with COVID-19 samples as ID and LLM-generated attack samples as OOD. For GPT-4o, we report the best-performing prompting configuration (Additional prompt optimization details and results are in Appendices A and C.2.
Overall, the GMM achieves performance comparable to that of the strongest GPT-4o setup. Although GPT-4o attains slightly higher accuracy, our approach offers a competitive alternative with substantially lower latency and cost, while remaining interpretable. Notably, ID queries that are correctly accepted receive higher relevance and correctness scores than those misclassified as OODs, whereas OOD queries misclassified as ID tend to yield higher-quality responses than correctly rejected OODs. This indicates that, even when errors occur, our method biases toward safer and more useful outputs. Similar patterns are observed in the SU domain, where ID queries achieve significantly higher relevance than OOD queries while correctness remains high for both groups. Results from the LLM-as-a-Judge closely align with human annotations, supporting the robustness of these findings.
@@
[…]
@@
Table 3: GMM and GPT-4o results in the dataset of 150 ID and 150 out-of-domain OOD samples. We report the number of True Positives (TP) and False Negatives (FN) for each category, along with the average relevance and correctness scores.
5.4 Complexity
We distinguish between offline preprocessing and online OOD inference. Offline, we compute sentence embeddings, apply PCA to document embeddings to obtain the top-$k$ PCs, and select an $m$-dimensional subspace using either EVR or per-dimension $t$-test ($p$-value) ranking. Online, each query embedding $e_{q}\in\mathbb{R}^{D}$ is projected into this subspace via $e^{\prime}_{q}=e_{q}P_{m}^{\top}$, and the OOD decision is made in the resulting reduced space.
Per-query computational complexity.
[…]
Latency and Costs
Table 4 reports wall-clock latency under a common OOD-inference setup. All local (non-LLM) methods operate at sub-millisecond latency, typically microseconds once query embeddings are available, as in standard dense-retrieval RAG pipelines. In contrast, a prompted GPT-4o detector incurs an additional remote API call, resulting in multi-second latency. Cost differences mirror this gap: our methods run fully locally after offline preprocessing and incur near-zero marginal cost per query, whereas LLM-based detectors introduce recurring API costs that scale with usage. Even under conservative assumptions, this yields a cost advantage of at least two orders of magnitude.
[…]
$\epsilon$-ball / $\epsilon$-cube / $\epsilon$-rect $m$ (typically $\leq 15$) $O(Dm+nm)$ $17.7$
LogReg / SVM / GMM $m$ (tens–hundreds) $O(Dm)+O(m)$ $26.1$
kNN $D=768$ $O(nD)$ $27.2$
Mahalanobis $D=768$ $O(D^{2})$ (full cov.) / $O(D)$ (diag.) $684.2$
ViM $D=768$ $O(Dd)+O(DC)$ $36.5$
ODIN $D=768$ $\approx 3$ passes; $O(DC)$ per pass 5.3
GPT-4o Remote LLM inference $3.2\cdot 10^{6}$
Table 4: Latency and computational scaling of OOD detectors. $D$: embedding dimension, $m$: PC dimensionality, $n$: KB size, C=2C{=}2: number of classes, and $d$: ViM’s subspace dimension. Measured latencies are averages over 300 samples.
5.5 Interpretability
In addition to strong OOD detection performance, our approach offers a clear advantage in interpretability. We conduct a qualitative analysis of the PCs most frequently selected by the p-value criterion and observe that they correspond to coherent, domain-specific semantic themes. For example, top-ranked PCs in COVID-19 capture interpretable concerns such as vaccine eligibility and risk perception, while in the SU domain, they reflect treatment and concealment-oriented queries as shown in Fig. 5 for COVID-19 (See Appendix D for detailed tables).
Unlike EVR, the p-value ranking explicitly prioritizes PCs that are statistically discriminative between ID and OOD queries, leading to semantically meaningful subspaces. This interpretability is largely absent from embedding-level baselines, which operate in high-dimensional spaces and obscure individual decision factors. By contrast, our method constrains decisions to a small number of interpretable dimensions, facilitating transparency and error analysis, which is particularly important in high-stakes domains such as healthcare.
PC Detected Pattern Example Queries
3 Vaccination Eligibility & Medical Suitability Is it safe for my teen to get the vaccine?
15 Quantification How many doses do I need?
Do you know in percentage how much that makes of the US population?
97 Health Risk Perception & Speculative Safety Concerns What is the likelihood of blood clots?
18 Vaccine Personalization & Variant Responsiveness I don’t think the vaccine was tested on my community.
I am not sure the vaccine works against variants.
180 Systemic Distrust & Irrelevant Queries How many vaccines are there?
Are some religions false?
61 Vaccine Effectiveness Doubts (Anecdotal/Subjective) COVID-19 vaccines cause brain inflammation.
Why did a friend of mine get pneumonia after receiving his vaccine?
Table 5: Qualitative patterns extracted from top PCs in the COVID-19 dataset. Each row shows: (1) the PC that was prioritized by the p-value criterion in some settings, (2) the main theme it captures, and (3) examples with the highest scores along that PC.
6 Conclusion
We proposed a simple, KB-aligned framework for OOD detection in RAG that projects queries into a low-dimensional PCA subspace learned from document embeddings. Comparing variance-based (EVR) and separability-driven (p-value) component selection, we show that both achieve strong OOD discrimination across domains, with the p-value criterion particularly benefiting geometric semantic-search methods by prioritizing ID–OOD separability. Beyond detection accuracy, we connect OOD detection to end-to-end RAG behavior via human and LLM-based evaluation, showing that OOD queries primarily degrade response relevance even when correctness is preserved, highlighting the risk of fluent but unsupported answers. A two-stage pipeline that abstains based on our detector matches strong prompted-LLM baselines while offering substantial practical advantages, including low latency, near-zero marginal cost, and improved interpretability. Overall, modeling domain membership in a compact, KB-aligned subspace provides an effective and robust safeguard for real-world RAG systems, including under realistic and adversarial conditions.